NvidiaIntely (nvidiaintely):

Determine the relationship between the lines that pass through the given points: Line A: (-3, 5) and (0, 7) Line B: (6, 2) and (9, 4) they are perpendicular they are parallel they are the same line they have no relations

NvidiaIntely (nvidiaintely):

@Kevin

OpenStudy (jiteshmeghwal9):

Let's form the line equations \((y-y_1)=\frac{(y_2-y_1)}{(x_2-x_1)}(x-x_1)\) Line A :- \((y-5)=\frac{7-5}{0-(-3)}(x-(-3))\) \((y-5)= \frac{2}{3} (x+3)\) Line B :- \((y-2)=\frac{4-2}{9-6}(x-6)\) \((y-2)=\frac{2}{3}(x-6)\)

OpenStudy (saksh1998):

calculate the slope and u will find your answer slopes are equal means lines are parallel

OpenStudy (jiteshmeghwal9):

As u can see that the slopes are equal so the lines are parallel besides there was no need to make line equation :P

OpenStudy (saksh1998):

with all due respect there is no use of forming equations please fan and medal

NvidiaIntely (nvidiaintely):

@Kevin